OpenStudy (anonymous):

I need help with these problems √72x^2 √250h^4k^5 √14q x 2√4q √63x^15y^9/7xy^11

OpenStudy (lgbasallote):

\[\huge \sqrt{75x ^{2}} or \sqrt{75} x ^{2}?\]

OpenStudy (diyadiya):

@lgbasallote 72 :D

OpenStudy (lgbasallote):

same thing :P

OpenStudy (diyadiya):

\[\sqrt{72 x^2}=\sqrt{72}*\sqrt{x^2}\]\[\sqrt{x^2}=x\]\[\sqrt{72}=6\sqrt{2}\]

OpenStudy (lgbasallote):

\[\large \sqrt{72x ^{2}} = \sqrt{(x)(x)(2)(2)(2)(3)(3)}\] factor out of the sqrt all that appeared twice

OpenStudy (anonymous):

how do I factor them out?

OpenStudy (diyadiya):

Just take out the numbers or variables that appeared twice from the square root =D

OpenStudy (anonymous):

sooo it would be √(x)(2)(3)?

OpenStudy (diyadiya):

Nope